# Mechanisms of electrostatic interactions between two charged dielectric spheres inside a polarizable medium: An effective-dipole analysis

### v1

@@ III Results and Discussion
A typical situation is the placement of two charged dielectric spheres in vacuum. In such a scenario, the effective dipole in the strongly polarized sphere 1 is all the way aligned along the direction of the Coulomb field from sphere 2. If sphere 2 carries a positive free charge, as in scenario 1(a) described below for two positively charged spheres, the induced negative charge in sphere 1 would be located closer to sphere 2 than its positive counterpart in the same sphere. If sphere 2 carries a negative free charge, as in scenario 1(b) described below for a positively charged sphere 1 and a negatively charged sphere 2, the induced positive charge in sphere 1 would be located closer to sphere 2 than its negative counterpart in the same sphere.
@@
Scenario 1(a): Like-charge repulsion or attraction
@@
Figure 2: Scenario 1(a): Like-charge repulsion or attraction for spheres immersed in a less polarizable medium at $k_{1}/k_{0}=30$ and $k_{2}/k_{0}=4$. The rows from top to bottom document results at respectively four different charge ratios: (a) $q_{2}/q_{1}=2$, (b) $q_{2}/q_{1}=4$, (c) $q_{2}/q_{1}=7$, and (d) $q_{2}/q_{1}=26$. The columns from left to right document results for (1) $F_{2}/\left|F_{\text{C}}\right|$ ($F_{2}$ is the electrostatic force on sphere 2 and $F_{\text{C}}$ is the Coulomb force), (2) $\widetilde{q}_{i}/q_{1}$ (relative magnitude of induced charge in each sphere with respect to $q_{1}$), (3) $(\widetilde{z}_{2}^{\pm}-z_{2})/a$ (relative positions of positive and negative induced charge in sphere 2 with respect to the sphere’s centre) and (4) a pictorial representation of positional order of induced charge and free charge in each sphere. In columns 1 to 3, the horizontal axis is set as $h/a$, representing a relative centre-to-centre separation with respect to the radius of the equal-sized spheres. In column 4, any induced charge is represented by a symbol with a dashed boundary, and any free charge is represented by a symbol with a solid boundary.
@@
Figure 3: Scenario 1(b): Opposite-charge attraction for spheres immersed in a less polarizable medium at $k_{1}/k_{0}=30$ and $k_{2}/k_{0}=4$. The rows from top to bottom document results at respectively four different charge ratios: (a) $q_{2}/q_{1}=-2$, (b) $q_{2}/q_{1}=-4$, (c) $q_{2}/q_{1}=-7$, and (d) $q_{2}/q_{1}=-26$. The columns from left to right document results for (1) $F_{2}/\left|F_{\text{C}}\right|$ ($F_{2}$ is the electrostatic force on sphere 2 and $F_{\text{C}}$ is the Coulomb force), (2) $\widetilde{q}_{i}/q_{1}$ (relative magnitude of induced charge in each sphere with respect to $q_{1}$), (3) $(\widetilde{z}_{2}^{\pm}-z_{2})/a$ (relative positions of positive and negative induced charge in sphere 2 with respect to the sphere’s centre) and (4) a pictorial representation of positional order of induced charge and free charge in each sphere. In columns 1 to 3, the horizontal axis is set as $h/a$, representing a relative centre-to-centre separation with respect to the radius of the equal-sized spheres. In column 4, any induced charge is represented by a symbol with a dashed boundary, and any free charge is represented by a symbol with a solid boundary.
@@
Figure 4: Plot of the reduced polarization force $(F_{2}-F_{\mathrm{C}})/\left|F_{\mathrm{C}}\right|$ against the reduced interparticle separation $h/a$ for cases of like-charge interactions as well as cases of opposite-charge interactions, at $k_{1}/k_{0}=30$ and $k_{2}/k_{0}=4$. The sub-figures (a) to (d) document results at $q_{2}/q_{1}=\pm 2$, $q_{2}/q_{1}=\pm 4$, $q_{2}/q_{1}=\pm 7$ and $q_{2}/q_{1}=\pm 26$, respectively. In all cases, the polarization force is attractive, be it for cases of like-charge interactions or opposite-charge interactions.
@@
Without loss of generality, for scenario 1(a) we consider like-charge interactions for respectively the charge ratios (a) $q_{2}/q_{1}=2$, (b) $q_{2}/q_{1}=4$, (c) $q_{2}/q_{1}=7$, and (d) $q_{2}/q_{1}=26$, at the dielectric-constant ratios $k_{1}/k_{0}=30$ and $k_{2}/k_{0}=4$. The results are summarized in Fig. 2. As shown in the first column, for decreasing interparticle separation, there is generally an increasing deviation of the electrostatic force $F_{2}$ on sphere 2 from the Coulomb force $F_{\text{C}}$, due to a strengthening of polarization-induced attraction. At sufficiently large absolute values of $q_{2}/q_{1}$ (e.g. $q_{2}/q_{1}=4$, $q_{2}/q_{1}=7$ and $q_{2}/q_{1}=26$), the Coulomb repulsion is overtaken by the polarization-induced attractive force at short interparticle separations such that the overall interparticle interaction becomes counterintuitively attractive. As shown in the second column, for decreasing interparticle separation, the magnitude of induced charge in each sphere as relative to $q_{1}$, i.e. $\widetilde{q}_{i}/q_{1}$, generally increases. For the four different values of $q_{2}/q_{1}$, we observed qualitative differences in the variation of $\widetilde{q}_{i}/q_{1}$ for decreasing interparticle separation: (a) At $q_{2}/q_{1}=2$, neither the value of $\widetilde{q}_{1}/q_{1}$ nor that of $\widetilde{q}_{2}/q_{1}$ exceeds unity, and the interparticle interaction remains repulsive all the way; (b) At $q_{2}/q_{1}=4$, the magnitudes of induced charge in both spheres are still all the way smaller than $q_{1}$, yet the value of $\widetilde{q}_{1}/q_{1}$ is not far from unity at short interparticle separations. This corresponds to a sufficiently strong effect of polarization for the occurrence of like-charge attraction; (c) At $q_{2}/q_{1}=7$, the value of $\widetilde{q}_{1}/q_{1}$ exceeds unity at small interparticle separations but the value of $\widetilde{q}_{2}/q_{1}$ is still all the way smaller than unity, and like-charge attraction occurs at short interparticle separations; (d) At $q_{2}/q_{1}=26$, both the value of $\widetilde{q}_{1}/q_{1}$ and that of $\widetilde{q}_{2}/q_{1}$ exceed unity at certain interparticle separations, and like-charge attraction occurs at short interparticle separations. As shown in the third column, the positions of induced positive charge and induced negative charge in sphere 2 as relative to the sphere’s centre exhibit a non-trivial variation for decreasing interparticle separation, at all four values of $q_{2}/q_{1}$. The results are translated into a pictorial representation in the fourth column, which illustrates the positional order of induced charge and free charge in each sphere. While the positional order of charge in sphere 1 is the same for any interparticle separation or charge ratio, i.e. with the induced positive charge on the left, the free positive charge in the middle, and the induced negative charge on the right, the positional order of charge in sphere 2 exhibits an interesting dependence on the interparticle separation and on the charge ratio, as discussed below:
@@
Scenario 1(b): Opposite-charge attraction
To investigate how two oppositely-charged particles in a less polarizable medium interact and facilitate a comparison with scenario 1(a), we have repeated the numerical investigations described above with a sign inversion of $q_{2}$ and with all other parameters remaining unchanged. The charge ratios considered are therefore (a) $q_{2}/q_{1}=-2$, (b) $q_{2}/q_{1}=-4$, (c) $q_{2}/q_{1}=-7$, and (d) $q_{2}/q_{1}=-26$, respectively. The results are summarized in Fig. 3. As shown in the first column, the electrostatic force between the two spheres remains uniformly attractive, regardless of the charge ratio nor the interparticle separation. Such attraction, which diverges from the conventional attraction between opposite charges as described by Coulomb’s law, is generally an enhancement of the Coulomb attraction. The larger the absolute value of the charge ratio, the stronger the enhancement. For decreasing interparticle separation, the electrostatic force $F_{2}$ on sphere 2 generally becomes more attractive and exhibits a greater deviation from the Coulomb force $F_{\text{C}}$, due to a strengthening of polarization-induced attraction. The fact that both the Coulomb force and the polarization-induced interaction are attractive in nature, as illustrated in Fig. 4 for the latter, excludes any possibility of opposite-charge repulsion. As shown in the second column of Fig. 3, for decreasing interparticle separation, the magnitude of induced charge in each sphere as relative to $q_{1}$, i.e. $\widetilde{q}_{i}/q_{1}$, generally increases, with features similar to scenario 1(a) for the four different values of $q_{2}/q_{1}$: (a) At $q_{2}/q_{1}=-2$, neither the value of $\widetilde{q}_{1}/q_{1}$ nor that of $\widetilde{q}_{2}/q_{1}$ exceeds unity; (b) At $q_{2}/q_{1}=-4$, the magnitudes of induced charge in both spheres are still all the way smaller than $q_{1}$, yet the value of $\widetilde{q}_{1}/q_{1}$ is not far from unity at short interparticle separations; (c) At $q_{2}/q_{1}=-7$, the value of $\widetilde{q}_{1}/q_{1}$ exceeds unity at small interparticle separations but the value of $\widetilde{q}_{2}/q_{1}$ is still all the way smaller than unity; (d) At $q_{2}/q_{1}=-26$, both the value of $\widetilde{q}_{1}/q_{1}$ and that of $\widetilde{q}_{2}/q_{1}$ exceed unity at certain interparticle separations.
As in the above analysis for scenario 1(a), the third column of Fig. 3 documents results on the positional order of induced charge and free charge in sphere 2, and the results are translated into a pictorial representation in the neighbouring fourth column. Regardless of the charge ratio nor the interparticle separation, the positional order of charge in either sphere is fixed, i.e. with the induced negative charge on the left, the free positive or negative charge in the middle, and the induced positive charge on the right. This is due to the following reason: For decreasing interparticle separation at any given charge ratio, the enhancement in the induced positive charge $\widetilde{q}_{1}^{+}$ in sphere 1 results in a further stabilization of the leftmost position of the induced negative charge $\widetilde{q}_{2}^{-}$ and the rightmost position of the induced positive charge $\widetilde{q}_{2}^{+}$ in the pictorial representation of sphere 2.
@@
Scenario 2: A more polarizable medium for both spheres
@@
A typical situation is the placement of two charged dielectric spheres inside a superdielectric medium Fromille and Phillips (2014), which has a dielectric constant of over $10^{5}$ at low frequencies ($<10^{-2}$z). In such a strongly polarizable medium, the region surrounding the strongly polarizing sphere 2 is filled with induced charge. As a result of significant influence from such induced charge, the effective dipole in sphere 1 is all the way aligned in a direction opposite to the Coulomb field from sphere 2. If sphere 2 carries a positive free charge, as in scenario 2(a) described below for two positively charged spheres, the induced positive charge in sphere 1 would be located closer to sphere 2 than its negative counterpart in the same sphere. If sphere 2 carries a negative free charge, as in scenario 2(b) described below for a positively charged sphere 1 and a negatively charged sphere 2, the induced negative charge in sphere 1 would be located closer to sphere 2 than its positive counterpart in the same sphere.
Scenario 2(a): Like-charge repulsion
@@
Figure 5: Scenario 2(a): Like-charge repulsion for spheres immersed in a more polarizable medium at $k_{1}/k_{0}=0.001$ and $k_{2}/k_{0}=0.08$. The rows from top to bottom document results at respectively four different charge ratios: (a) $q_{2}/q_{1}=3$, (b) $q_{2}/q_{1}=8$, (c) $q_{2}/q_{1}=16$, and (d) $q_{2}/q_{1}=80$. The columns from left to right document results for (1) $F_{2}/\left|F_{\text{C}}\right|$ ($F_{2}$ is the electrostatic force on sphere 2 and $F_{\text{C}}$ is the Coulomb force), (2) $\widetilde{q}_{i}/q_{1}$ (relative magnitude of induced charge in each sphere with respect to $q_{1}$), (3) $(\widetilde{z}_{2}^{\pm}-z_{2})/a$ (relative positions of positive and negative induced charge in sphere 2 with respect to the sphere’s centre) and (4) a pictorial representation of positional order of induced charge and free charge in each sphere. In columns 1 to 3, the horizontal axis is set as $h/a$, representing a relative centre-to-centre separation with respect to the radius of the equal-sized spheres. In column 4, any induced charge is represented by a symbol with a dashed boundary, and any free charge is represented by a symbol with a solid boundary.
@@
Figure 6: Plot of the reduced polarization force $(F_{2}-F_{\mathrm{C}})/\left|F_{\mathrm{C}}\right|$ against the reduced interparticle separation $h/a$ for cases of like-charge interactions as well as cases of opposite-charge interactions, at $k_{1}/k_{0}=0.001$ and $k_{2}/k_{0}=0.08$. The sub-figures (a) to (d) document results at $q_{2}/q_{1}=\pm 3$, $q_{2}/q_{1}=\pm 8$, $q_{2}/q_{1}=\pm 16$ and $q_{2}/q_{1}=\pm 80$, respectively. In all cases, the polarization force is repulsive, be it for cases of like-charge interactions or opposite-charge interactions.
For cases of two like-charged particles immersed in a more polarizable medium with $k_{1}/k_{0}=0.001$ and $k_{2}/k_{0}=0.08$, we have conducted numerical investigations for four different charge ratios at (a) $q_{2}/q_{1}=3$, (b) $q_{2}/q_{1}=8$, (c) $q_{2}/q_{1}=16$, and (d) $q_{2}/q_{1}=80$, respectively. The results are summarized in Fig. 5. As shown in the first column, the electrostatic force between the two spheres remains uniformly repulsive, regardless of the charge ratio nor the interparticle separation. Such repulsion, which diverges from the conventional repulsion between like charges as described by Coulomb’s law, is generally an enhancement of the Coulomb repulsion. The larger the absolute value of the charge ratio, the stronger the enhancement. For decreasing interparticle separation, the electrostatic force $F_{2}$ on sphere 2 generally becomes more repulsive and exhibits a greater deviation from the Coulomb force $F_{\text{C}}$, due to a strengthening of polarization-induced repulsion. The fact that both the Coulomb force and the polarization-induced interaction are repulsive in nature, as illustrated in Fig. 6 for the latter, excludes any possibility of like-charge attraction. As shown in the second column of Fig. 5, for decreasing interparticle separation, the magnitude of induced charge in each sphere as relative to $q_{1}$, i.e. $\widetilde{q}_{i}/q_{1}$, generally increases, with features similar to scenarios 1(a) and 1(b) for the four different values of $q_{2}/q_{1}$: (a) At $q_{2}/q_{1}=3$, neither the value of $\widetilde{q}_{1}/q_{1}$ nor that of $\widetilde{q}_{2}/q_{1}$ exceeds unity; (b) At $q_{2}/q_{1}=8$, the magnitudes of induced charge in both spheres are still all the way smaller than $q_{1}$, yet the value of $\widetilde{q}_{1}/q_{1}$ is not far from unity at short interparticle separations; (c) At $q_{2}/q_{1}=16$, the value of $\widetilde{q}_{1}/q_{1}$ exceeds unity at small interparticle separations but the value of $\widetilde{q}_{2}/q_{1}$ is still all the way smaller than unity; (d) At $q_{2}/q_{1}=80$, both the value of $\widetilde{q}_{1}/q_{1}$ and that of $\widetilde{q}_{2}/q_{1}$ exceed unity at certain interparticle separations.
@@
As in the preceding analysis, the third column of Fig. 5 documents results on the positional order of induced charge and free charge in sphere 2, and the results are translated into a pictorial representation in the neighbouring fourth column. Regardless of the charge ratio nor the interparticle separation, the positional order of charge in sphere 1 is fixed, i.e. with the induced negative charge on the left, the free positive charge in the middle, and the induced positive charge on the right. This is due to the following reason: For decreasing separation at any given charge ratio, the enhancement in the induced negative charge in the medium region between the particles results in a further stabilization of the leftmost position of the induced negative charge $\widetilde{q}_{1}^{-}$ and the rightmost position of the induced positive charge $\widetilde{q}_{1}^{+}$ in the pictorial representation of sphere 1.
@@
Scenario 2(b): Opposite-charge attraction or repulsion
Figure 7: Scenario 2(b): Opposite-charge attraction or repulsion for spheres immersed in a more polarizable medium at $k_{1}/k_{0}=0.001$ and $k_{2}/k_{0}=0.08$. The rows from top to bottom document results at respectively four different charge ratios: (a) $q_{2}/q_{1}=-3$, (b) $q_{2}/q_{1}=-8$, (c) $q_{2}/q_{1}=-16$, and (d) $q_{2}/q_{1}=-80$. The columns from left to right document results for (1) $F_{2}/\left|F_{\text{C}}\right|$ ($F_{2}$ is the electrostatic force on sphere 2 and $F_{\text{C}}$ is the Coulomb force), (2) $\widetilde{q}_{i}/q_{1}$ (relative magnitude of induced charge in each sphere with respect to $q_{1}$), (3) $(\widetilde{z}_{2}^{\pm}-z_{2})/a$ (relative positions of positive and negative induced charge in sphere 2 with respect to the sphere’s centre) and (4) a pictorial representation of positional order of induced charge and free charge in each sphere. In columns 1 to 3, the horizontal axis is set as $h/a$, representing a relative centre-to-centre separation with respect to the radius of the equal-sized spheres. In column 4, any induced charge is represented by a symbol with a dashed boundary, and any free charge is represented by a symbol with a solid boundary.

### v2

@@ IV Mechanisms of interparticle interactions
If two charged spheres are placed in a less polarizable medium, for example in a vacuum, the effective dipole in the strongly polarized sphere 1 must align along the direction of the strong Coulomb field from sphere 2. If sphere 2 is positively charged, as in scenario B described below for two positively charged spheres, the induced negative charge in sphere 1 would be located closer to sphere 2 than its positive counterpart in the same sphere. Conversely, if sphere 2 is negatively charged, as in scenario D described below for a positively charged sphere 1 and a negatively charged sphere 2, the induced positive charge in sphere 1 would be located closer to sphere 2 than its negative counterpart in the same sphere. Scenarios B and D could be realized experimentally in a vacuum ($k_{0}=1$) with a lanthana-based sphere 1 ($k_{1}\sim 30$) and a silica-based sphere 2 ($k_{2}\sim 4$).
@@
Scenario B: Like-charge repulsion or attraction in a less polarizable medium
@@
Figure 5: Scenario B: Like-charge repulsion or attraction for spheres immersed in a less polarizable medium at $k_{1}/k_{0}=30$ and $k_{2}/k_{0}=4$. The rows from top to bottom document results at respectively four different charge ratios: (a) $q_{2}/q_{1}=2$, (b) $q_{2}/q_{1}=4$, (c) $q_{2}/q_{1}=7$, and (d) $q_{2}/q_{1}=26$. The columns from left to right document results for (1) $F_{2}/\left|F_{\text{C}}\right|$ ($F_{2}$ is the electrostatic force on sphere 2 and $F_{\text{C}}$ is the Coulomb force), (2) $\widetilde{q}_{i}/q_{1}$ (relative magnitude of induced charge in each sphere with respect to $q_{1}$), (3) $(\widetilde{z}_{2}^{\pm}-z_{2})/a$ (relative positions of positive and negative induced charge in sphere 2 with respect to the sphere’s centre) and (4) a pictorial representation of positional order of induced charge and free charge in each sphere. In columns 1 to 3, the horizontal axis is set as $h/a$, representing a relative centre-to-centre separation with respect to the radius of the equal-sized spheres. In column 4, any induced charge is represented by a symbol with a dashed boundary, and any free charge is represented by a symbol with a solid boundary.
@@
Figure 6: Scenario D: Opposite-charge attraction for spheres immersed in a less polarizable medium at $k_{1}/k_{0}=30$ and $k_{2}/k_{0}=4$. The rows from top to bottom document results at respectively four different charge ratios: (a) $q_{2}/q_{1}=-2$, (b) $q_{2}/q_{1}=-4$, (c) $q_{2}/q_{1}=-7$, and (d) $q_{2}/q_{1}=-26$. The columns from left to right document results for (1) $F_{2}/\left|F_{\text{C}}\right|$ ($F_{2}$ is the electrostatic force on sphere 2 and $F_{\text{C}}$ is the Coulomb force), (2) $\widetilde{q}_{i}/q_{1}$ (relative magnitude of induced charge in each sphere with respect to $q_{1}$), (3) $(\widetilde{z}_{2}^{\pm}-z_{2})/a$ (relative positions of positive and negative induced charge in sphere 2 with respect to the sphere’s centre) and (4) a pictorial representation of positional order of induced charge and free charge in each sphere. In columns 1 to 3, the horizontal axis is set as $h/a$, representing a relative centre-to-centre separation with respect to the radius of the equal-sized spheres. In column 4, any induced charge is represented by a symbol with a dashed boundary, and any free charge is represented by a symbol with a solid boundary.
@@
Figure 7: Plot of the rescaled polarization force $(F_{2}-F_{\mathrm{C}})/\left|F_{\mathrm{C}}\right|$ against the rescaled interparticle separation $h/a$ for cases of like-charge interactions as well as cases of opposite-charge interactions, at $k_{1}/k_{0}=30$ and $k_{2}/k_{0}=4$. The sub-figures (a) to (d) document results at $q_{2}/q_{1}=\pm 2$, $q_{2}/q_{1}=\pm 4$, $q_{2}/q_{1}=\pm 7$ and $q_{2}/q_{1}=\pm 26$, respectively. In all cases, the polarization force is attractive, be it for cases of like-charge interactions or opposite-charge interactions.
@@
For scenario B we consider like-charge interactions for respectively the charge ratios (a) $q_{2}/q_{1}=2$, (b) $q_{2}/q_{1}=4$, (c) $q_{2}/q_{1}=7$, and (d) $q_{2}/q_{1}=26$ with $q_{1}=1\text{e}$, at the dielectric-constant ratios $k_{1}/k_{0}=30$ and $k_{2}/k_{0}=4$ with $k_{0}=1$. The results are summarized in Fig. 5. As shown in the first column, for decreasing interparticle separation, there is generally an increasing deviation of the electrostatic force $F_{2}$ on sphere 2 from the Coulomb force $F_{\text{C}}$, due to a strengthening of polarization-induced attraction. At sufficiently large absolute values of $q_{2}/q_{1}$ (e.g. $q_{2}/q_{1}=4$, $q_{2}/q_{1}=7$ and $q_{2}/q_{1}=26$), the Coulomb repulsion is overtaken by the polarization-induced attractive force at short interparticle separations such that the overall interparticle interaction becomes counterintuitively attractive.
As shown in the second column, for decreasing interparticle separation, the magnitude of induced charge in each sphere as relative to $q_{1}$, i.e. $\widetilde{q}_{i}/q_{1}$, generally increases. For the four different values of $q_{2}/q_{1}$, we observed qualitative differences in the variation of $\widetilde{q}_{i}/q_{1}$ for decreasing interparticle separation: (a) At $q_{2}/q_{1}=2$, neither the value of $\widetilde{q}_{1}/q_{1}$ nor that of $\widetilde{q}_{2}/q_{1}$ exceeds unity, and the interparticle interaction remains repulsive all the way; (b) At $q_{2}/q_{1}=4$, the magnitudes of induced charge in both spheres are still all the way smaller than $q_{1}$, yet the value of $\widetilde{q}_{1}/q_{1}$ is not far from unity at short interparticle separations. This corresponds to a sufficiently strong effect of polarization for the occurrence of like-charge attraction; (c) At $q_{2}/q_{1}=7$, the value of $\widetilde{q}_{1}/q_{1}$ exceeds unity at small interparticle separations but the value of $\widetilde{q}_{2}/q_{1}$ is still all the way smaller than unity, and like-charge attraction occurs at short interparticle separations; (d) At $q_{2}/q_{1}=26$, both the value of $\widetilde{q}_{1}/q_{1}$ and that of $\widetilde{q}_{2}/q_{1}$ exceed unity at certain interparticle separations, and like-charge attraction occurs at short interparticle separations.
As shown in the third column, the positions of induced positive charge and induced negative charge in sphere 2 as relative to the sphere’s centre exhibit a non-trivial variation for decreasing interparticle separation, at all four values of $q_{2}/q_{1}$. The results are translated into a pictorial representation in the fourth column, which illustrates the positional order of induced charge and free charge in each sphere. While the positional order of charge in sphere 1 is the same for any interparticle separation or charge ratio, i.e. with the induced positive charge on the left, the free positive charge in the middle, and the induced negative charge on the right, the positional order of charge in sphere 2 exhibits an interesting dependence on the interparticle separation and on the charge ratio, as discussed below:
@@
Scenario D: Opposite-charge attraction in a less polarizable medium
To investigate how two oppositely-charged particles in a less polarizable medium interact and facilitate a comparison with scenario B, we have repeated the numerical investigations described above with a sign inversion of $q_{2}$ and with all other parameters remaining unchanged. The charge ratios considered are therefore (a) $q_{2}/q_{1}=-2$, (b) $q_{2}/q_{1}=-4$, (c) $q_{2}/q_{1}=-7$, and (d) $q_{2}/q_{1}=-26$, respectively, with $q_{1}=1\text{e}$. The results are summarized in Fig. 6. As shown in the first column, the electrostatic force between the two spheres is uniformly attractive, regardless of the charge ratio nor the interparticle separation. Such attraction, which diverges from the conventional attraction between opposite charges as described by Coulomb’s law, is generally an enhancement of the Coulomb attraction. The larger the absolute value of the charge ratio, the stronger the enhancement. For decreasing interparticle separation, the electrostatic force $F_{2}$ on sphere 2 generally becomes more attractive and exhibits a greater deviation from the Coulomb force $F_{\text{C}}$, due to a strengthening of polarization-induced attraction. The fact that both the Coulomb force and the polarization-induced interaction are attractive in nature, as illustrated in Fig. 7 for the latter, excludes any possibility of opposite-charge repulsion.
Regardless of the charge ratio nor the interparticle separation, the positional order of charge in either sphere is fixed, i.e. with the induced negative charge on the left, the free positive or negative charge in the middle, and the induced positive charge on the right. This is because, for decreasing interparticle separation at any given charge ratio, the enhancement in the induced positive charge $\widetilde{q}_{1}^{+}$ in sphere 1 results in a further stabilization of the leftmost position of the induced negative charge $\widetilde{q}_{2}^{-}$ and the rightmost position of the induced positive charge $\widetilde{q}_{2}^{+}$ in the pictorial representation of sphere 2.
@@
Cases of a more polarizable medium
@@
A possible experimental realization is the placement of a non-polarizable sphere 1 ($k_{1}\sim 1$) and a water-based sphere 2 ($k_{2}\sim 80$), inside a superdielectric medium Fromille and Phillips (2014) of dielectric constant $k_{0}=1000$. In such a strongly polarizable medium, the region surrounding the strongly polarizing sphere 2 is filled with induced charge. As a result of significant influence from such induced charge, the effective dipole in sphere 1 must align in a direction opposite to the strong Coulomb field from sphere 2.
Scenario A: Like-charge repulsion in a more polarizable medium
@@
Figure 8: Scenario A: Like-charge repulsion for spheres immersed in a more polarizable medium at $k_{1}/k_{0}=0.001$ and $k_{2}/k_{0}=0.08$. The rows from top to bottom document results at respectively four different charge ratios: (a) $q_{2}/q_{1}=3$, (b) $q_{2}/q_{1}=8$, (c) $q_{2}/q_{1}=16$, and (d) $q_{2}/q_{1}=80$. The columns from left to right document results for (1) $F_{2}/\left|F_{\text{C}}\right|$ ($F_{2}$ is the electrostatic force on sphere 2 and $F_{\text{C}}$ is the Coulomb force), (2) $\widetilde{q}_{i}/q_{1}$ (relative magnitude of induced charge in each sphere with respect to $q_{1}$), (3) $(\widetilde{z}_{2}^{\pm}-z_{2})/a$ (relative positions of positive and negative induced charge in sphere 2 with respect to the sphere’s centre) and (4) a pictorial representation of positional order of induced charge and free charge in each sphere. In columns 1 to 3, the horizontal axis is set as $h/a$, representing a relative centre-to-centre separation with respect to the radius of the equal-sized spheres. In column 4, any induced charge is represented by a symbol with a dashed boundary, and any free charge is represented by a symbol with a solid boundary.
Figure 9: Scenario C: Opposite-charge attraction or repulsion for spheres immersed in a more polarizable medium at $k_{1}/k_{0}=0.001$ and $k_{2}/k_{0}=0.08$. The rows from top to bottom document results at respectively four different charge ratios: (a) $q_{2}/q_{1}=-3$, (b) $q_{2}/q_{1}=-8$, (c) $q_{2}/q_{1}=-16$, and (d) $q_{2}/q_{1}=-80$. The columns from left to right document results for (1) $F_{2}/\left|F_{\text{C}}\right|$ ($F_{2}$ is the electrostatic force on sphere 2 and $F_{\text{C}}$ is the Coulomb force), (2) $\widetilde{q}_{i}/q_{1}$ (relative magnitude of induced charge in each sphere with respect to $q_{1}$), (3) $(\widetilde{z}_{2}^{\pm}-z_{2})/a$ (relative positions of positive and negative induced charge in sphere 2 with respect to the sphere’s centre) and (4) a pictorial representation of positional order of induced charge and free charge in each sphere. In columns 1 to 3, the horizontal axis is set as $h/a$, representing a relative centre-to-centre separation with respect to the radius of the equal-sized spheres. In column 4, any induced charge is represented by a symbol with a dashed boundary, and any free charge is represented by a symbol with a solid boundary.
@@
Figure 10: Plot of the rescaled polarization force $(F_{2}-F_{\mathrm{C}})/\left|F_{\mathrm{C}}\right|$ against the rescaled interparticle separation $h/a$ for cases of like-charge interactions as well as cases of opposite-charge interactions, at $k_{1}/k_{0}=0.001$ and $k_{2}/k_{0}=0.08$. The sub-figures (a) to (d) document results at $q_{2}/q_{1}=\pm 3$, $q_{2}/q_{1}=\pm 8$, $q_{2}/q_{1}=\pm 16$ and $q_{2}/q_{1}=\pm 80$, respectively. In all cases, the polarization force is repulsive, be it for cases of like-charge interactions or opposite-charge interactions.
We have conducted numerical investigations for four different charge ratios at (a) $q_{2}/q_{1}=3$, (b) $q_{2}/q_{1}=8$, (c) $q_{2}/q_{1}=16$, and (d) $q_{2}/q_{1}=80$, respectively, with $q_{1}=1\text{e}$. The results, which are summarized in Fig. 8, are analogous to scenario D, where in either scenario the corresponding Coulomb interaction is enhanced. As an enhancement of the Coulomb repulsion, the electrostatic force between the two spheres is uniformly repulsive, where the same repulsive nature of the Coulomb force and the polarization-induced interaction (Fig. 10) excludes any possibility of like-charge attraction.
@@
In this scenario, the positional order of charge in sphere 1 is fixed, i.e. with the induced negative charge on the left, the free positive charge in the middle, and the induced positive charge on the right. For decreasing separation at any given charge ratio, the enhancement in the induced negative charge in the medium region between the particles results in a further stabilization of the leftmost position of the induced negative charge $\widetilde{q}_{1}^{-}$ and the rightmost position of the induced positive charge $\widetilde{q}_{1}^{+}$ in the pictorial representation of sphere 1.
@@
Scenario C: Opposite-charge attraction or repulsion in a more polarizable medium